OpenStudy (cookie_2046):

A 10-foot ladder is leaning against a tree. The bottom of the ladder is 4 feet away from the bottom of the tree. Approximately how high up the tree does the top of the ladder reach? A. 2.4 feet B. 6 feet C. 9.2 feet D. 10.8 feet

OpenStudy (anonymous):

in this problem ladder is hypotunes c 10 ft distance frm tree b 4 ft and hieght of tree a x by pethagoras theoram c^2= a^2 +b^2 a^2 = c^2 - b^2 a^2 = 10^2 - 4^2